OpenStudy (lukecrayonz):

Write the equation of an ellipse with foci at (-4, 0) and (-4, -6) and major axis of 10

OpenStudy (lukecrayonz):

@dpaInc yeahhhhhhhhh

OpenStudy (lukecrayonz):

I got it!

OpenStudy (anonymous):

that's great...:)

OpenStudy (anonymous):

for the foci is at (-4, 0) and (-4, -6), the major axis is parallel to y-axis the a=10/2=5, c=(0-(-6))/2=3, so \[b=\sqrt{5^{2}-3^{2}}=4\] the basic equation should be and the foci of it is (0,3) & (0,-3) move them to (-4, 0) and (-4, -6) so we can have \[(y-3) ^{2}/25+(x+4) ^{2}/16=1\]

OpenStudy (anonymous):

 yep... @AndrewNJ got it...